In the market for apples in Country A, cost curves are shown for both monopoly and perfect competition. Under perfect competition 400 apples are produced at a price of $5 while if the market were operated by Apple Kings (A monopoly) only 150 apples would be produced at a price of $15 per apple. The y-intercept of both Marginal revenue and the demand curve is at $30.

(a) Calculate the consumer surplus if the market were operated by Apple Kings

(b) Identify and calculate the Consumer surplus under perfect competition

(c) What is the difference in consumer surplus if the market should move from perfect competition to monopoly and is this alignment with the theory?
Answer #1

Answer:

Consumer surplus = ½*(highest willingness to apy - price )*quantity

(a) In monopoly market, CS = ½*(30 - 15) * 150 = $1,125
(b) In perfec competition, CS = ½*(30 - 5) * 400 = $5,000
(c) If the market moves from perfect competition to monopoly, consumer surplus will decrease by $3,875 (5,000 - 1,125 = 3,875). This is in alignment with the theory because in monopoly market consumer surplus is lost because quantity is determined by the intersection of MC curve with MR curve, not demand curve, like in perfect competition. So there is welfare loss (deadweight loss) in monopoly market.